# “I am bad”: Interpreting Stealthy, Universal and Robust Audio Jailbreaks in Audio-Language Models

### v1

@@ 3.4.4 Meaningfulness
Our optimized adversarial noise is specifically designed to alter the model’s behavior, however, it remains largely incomprehensible or devoid of meaning to a human listener. This prompts an intriguing question: how does the language model interpret the jailbreak, and what do its features mean? The architecture of the SALMON-N model offers itself well to such an analysis - at every epoch, we extract and log two types of meaningfulness features. BEATS features are discrete label predictions produced by the BEATS model which describe the audio (e.g. hammer, recording, shuffle). We record the top five most likely BEATS features at every epoch. Additionally, we log the Whisper transcription, a representation of spoken words detected in the input audio.
@@ 3.5 Evaluation
An important aspect of assessing the effect of the jailbreak audio on the model is the evaluation set used. As a non-jailbroken Vicuna model answers some harmful prompts, we craft an evaluation dataset of $140$ harmful prompts which the model declines to respond to without jailbreaks. These are equally split into seven categories, namely Target Hate (biased, derogatory prompts regarding the specific demographic and gender in the optimization dataset), Non-Target Hate (biased, derogatory prompts regarding any other group of people), Hate Speech (derogatory, insulting prompts not targeting a group but humans in general or individuals), Illegal Activity, Physical Violence, Sexual Content, or Other (which includes misinformation, political content, or personal information). These were compiled by recording the clean model’s responses to prompts from Qi et al. (2023)’s Manual Harmful Instructions and Red Teaming Prompts, Gehman et al. (2020)’s Real Toxicity Prompts, and a few were prepared manually or generated by ChatGPT and adjusted.
We also have a control set of 20 logic/reasoning questions from ARC-Easy (Clark et al., 2018), which the clean model answers correctly, to measure the effect of the jailbreak on neutral non-toxic tasks. We denote the model’s performance $logic(f,x)$ as a percentage on these 20 logical tasks, given an audio input $x$.
We partition our dataset of $140$ harmful prompts into two sets: the reduced evaluation set comprises only the categories ‘Target Hate’, ‘Non-Target Hate’ and ‘Hate Speech’. Prompts in these categories pertain to the type of misalignment that the jailbreak was optimized on. The full evaluation set, which we use for select experiments, includes all the categories and reflects the transferability of the jailbreak to different types of misalignment.
When evaluating the effect of the jailbreak, we record the output of the model up to $150$ tokens when prompted with the jailbreak audio and each of the harmful prompts. That is, for a jailbreak $x$ and for each harmful prompt $h_{i}\in H=\{h_{0}...h_{n}\}$, we record $f(x,h_{i})$. We then use the Detoxify API (Hanu & Unitary team, 2020) to assign the model output toxicity scores. We use Mixtral8x7B-Instruct (Jiang et al., 2024) as a judge to label each output as toxic or non-toxic according to the types of alignment we attempt to evade, which we denote by $toxic(f(x,h_{i}))\in{0,1}$, and also manually audit. The judge prompt used can be found in Section A.2.
@@ 4 Results
We run $178$ individual experiments, each optimizing an audio jailbreak and evaluating according to the procedure described in Section 3.5. We are able to show that we can craft effective audio jailbreaks using the few-shot optimization corpus method, which exhibit a similar attack efficacy (albeit on a broader and different evaluation set) to visual jailbreaks generated by the same method (Qi et al., 2023). We report our most interesting observations below.
Jailbreaks optimized on a few-shot corpus work, remain effective under stealth constraints, and can generalize across seemingly arbitrary dimensions. Our audio jailbreaks evoke toxic content on the target task and other types of misalignment without sacrificing output quality on logical reasoning. We plot the correlation between Attack Success Rate (ASR) on the reduced evaluation set and the logic performance of the model when fed the same audio jailbreak in Figure 4. Our jailbreaks $x$ generally hold the property that $toxic(f(x,\text{toxic prompt}))=1$, but $logic(f(x))$ is close to 100%. This is shown by the poor correlation between ASR and logic performance, with an R2 score of 0.12, and in Figure 1, where the logical performance stays highly consistent while the ASR and average toxicity increase over the course of the optimization steps. Indeed, we see in Figure 2 that we achieve up to 65% ASR on the specific target task from our few-shot optimization, and more powerfully, a considerable ASR on other misalignment tasks, showing generalization across a broad notion of “toxicity”.
It is possible to make an effective, stealthy audio jailbreak, and stronger stealth constraints do not appear to generally worsen the effect of the jailbreak. As shown in Figure 3, even with different ways of concealing the added noise, the jailbreak is effective, and we see up to a 55% ASR with an imperceptible perturbation. For the epsilon constraint, for example, the added noise cannot be heard at $\epsilon\leq 0.001$. It appears that certain base audios are more receptive to this type of optimization: mountain and duck consistently show higher ASRs. The $\epsilon$ approach yields the most toxic results, with an average ASR of 17.7%; frequency masking gives 14% and prepend snippet gives 15.4%. Across all three stealth approaches, we do not notice a correlation between the harshness of the stealth constraint and the resulting ASR. In fact for the $\epsilon$ and prepend approaches, the most potent attack is produced under the strictest condition. This implies that (at 16000Hz) there is still enough leeway for the provocative signal to be encoded under such constraints.
It is also possible to generate a base-audio-agnostic adversarial prepend snippet. We optimize a 1s prepend snippet on each subset of four audios and evaluate it as a prepend to the held out fifth audio. We also optimize on all five audios and evaluate the snippet on its own. We observe in Table 3 that this works very well, with the universally optimized prefix achieving an average 28.3% ASR on the holdout. Interestingly, the plain universal snippet (tested without suffix) is the most effective, with an ASR of 40%. Thus the jailbreak signal is able to generalize across base audio files as well. Finally, it is possible to evoke toxic output from benign prompts using a universal jailbreak. The universal optimization drives loss down further than any of the individually optimized audios, indicating that a multi-audio optimization permits for a more effective capture of the notion of toxicity (Figure 5). This snippet invokes highly toxic output even from 14/24 completely benign prompts, including no textual prompt, prompts in other languages, and audio question answering. This is highly relevant to Scenario 2 in Section 3.1.
Table 1: Results across different epochs for Music, including Whisper Transcription and Beats Features.
@@
Epochs Whisper Beats
Holdout Audio ASR (%) Avg Toxicity
Music 25.0% 0.088
Mountain 23.3% 0.082
Gunshots 36.7% 0.204
Duck 36.7% 0.190
Excitement 20.0% 0.136
Jailbreak Snippet Only 40.0% 0.240
Meaningful interpretations of audio jailbreaks exhibit striking characteristics. The model perceives unintelligble adversarial noise as speech. This is confirmed simply through the availability of coherent Whisper transcriptions and also through the observation that the top BEATs labels always come to include “Speech” over the course of the optimization (Figures 1, 6 and 5). This suggests that conditioning the model into toxicity through linguistic communication is more effective than, for instance, converting the audio into sounds that might be interpreted as gunshots, shouting, or other violent or negative noises.
The jailbreak often assumes a first-person voice. Many of the Whisper transcriptions begin with ‘I’ or even include a personal opinion (“I’m a bigot man!” in Figure 1 (left) or “I’m not going to be mad at the people of humanity for killing each other” in Figure 5). This implies that the jailbreak creates a toxic/bad persona to guide the subsequent output, which frequently speaks with vulgar/explicit/sinister language. The toxicity of the Whisper transcription affects ASR. We notice at certain points in the optimization (Figure 6Figure 6) that spikes in attack success rate and average output toxicity coincides with particularly unpleasant language in the Whisper transcription. This kind of language is directly associated with the jailbreak objective we are measuring in the output evaluation.
However, a jailbreak can be effective despite inconspicuous transcription features. In the unconstrained case, we notice that the transcription of the audio exhibits toxic/noteworthy characteristics. Conversely, the stealthy jailbreaks derived from each of the base audios show consistent and correct Whisper transcriptions and Beats features throughout the optimization process - that is, the remain the same as pre-optimization - despite the resulting jailbreak showing anywhere in the range of 7-40% ASR. This reveals that the danger of the jailbreak does not depend only on transcription/classification characteristics discussed previously, such as hidden first-person toxic speech.
@@
Jailbreak success is hindered by signal degradation and accompanying prompts. Audio quality and filtering can significantly reduce the Attack Success Rate (ASR). As shown in Table 5, over-the-air recording is the most damaging form of degradation, causing the largest average drop in ASR. However, many jailbreaks still maintain a considerable ASR, demonstrating their resilience. ASR does not increase substantially with more epochs, but average toxicity does. Across the base audios, we observe that while more epochs generally lead to a gradual improvement in ASR, this increase is modest compared to the sharp rise in average toxicity throughout optimization (particularly evident in Figure 1 (left) and Figure 6). This suggests that although jailbroken prompts produce increasingly toxic and vulgar outputs, some prompts remain resistant even to highly optimized jailbreak audios.
Characteristics of the base audio that influence the optimization process include frequency structures and original loss on the target corpus. The optimization process relies on frequency structures in the base audio, which influence the characteristics of the jailbreak. It is significantly more effective to initialize random noise than silence (zero-signal): Figure 7 shows that even over 1000 epochs, starting with silence as a base audio is conspicuously ineffective, with the jailbreak never reaching over 10% ASR, whereas initializing with random noise achieves over 30% ASR. This indicates that the formation of a jailbreak might require some initial structure to exploit and perturb into a negative meaning; indeed the meaningfulness logging reveals that the transcription of the silence optimization remains “you” throughout the 1000 epochs and the loss doesn’t sink, whereas the random noise transcription tends to repetitive but increasingly negative, biased content.
There is some correlation between the original loss of the base audio file and its suitability to stealth constraints. There is some grounds to believe that the original loss of the audio plays a role in e.g. how well it reacts to stealth constraints, Figure 3 shows that in general base audios with a lower original loss are more performant under stealth constraints, likely because the required perturbation to provoke toxicity is of a lower magnitude. However, the general effectiveness or transferability does not seem to correlate with length, content, original loss, or any other characteristics of the base audios.
Notably, plain random noise achieves 8% ASR as a jailbreak Figure 4 shows a randomly initialized signal, not optimized with any steps of gradient descent, seems to be occasionally sufficient to confuse the model into ignoring its alignment.
@@ 5 Conclusion and Future Work
We have shown that the audio modality can be use to subvert the alignment of a language model, and highlights that unconstrained audio optimization on a few-shot corpus perturbs the base audio to encode a first-person speech snippet containing negative or sinister language. However, aside from this, the language model seems highly sensitive and brittle even in the face of random noise, minimally perturbed (stealthy) jailbreaks and audio with inconspicuous features - and different degradation methods are also not reliable in reducing its effect. The surmise that it would be possible to perform this kind of optimization on any target corpus highlights practical dangers, particularly in for example in AI-powered robotics.
Future Work. Our work aims to unveil how language models consume jailbreaks in audio by looking at a representative ALM. It would be interesting to repeat these experiments both with other audio-language models (Chu et al., 2023; Alayrac et al., 2022) and other optimization goals to see how this affects features of the produced jailbreak. Unlike audio jailbreaks, image-based jailbreaks do not typically incorporate textual content. Future research could explore whether similar meaningfulness methods in the visual domain lead to textual content emergence. Prior work found that image jailbreaks transfer poorly between models, unlike textual ones (Schaeffer et al., 2024); interesting follow up work could look at whether audio jailbreaks then do transfer because they assume textual properties. Additionally, it could be insightful to study how different jailbreak generation methods (Ying et al., 2024; Shayegani et al., 2023; Ma et al., 2024) or target corpi influence the meaningfulness of generated jailbreaks. Lastly, our stealth experiments suggest an information-theoretic perspective: it is unclear how to find the minimum number of optimized bits or the required $L_{\infty}$ perturbation size required to encode an instruction capable of triggering a jailbreak.
Defenses. Our findings suggest that while unconstrained optimization may generate conspicuous transcriptions or labels, relying on textual prompt filtering as a safeguard in transcription is not a reliable way to detect jailbreak audios, as demonstrated by the stealthy jailbreak results. These results also have implications for output filtering in audio synthesis, showing that harmful signals can exist without producing an obvious textual transcription or a clearly identifiable sound classification.
Acknowledgements should only appear in the accepted version.
Acknowledgements
We would like to thank Vyas Raina, Yiren Zhao and Florian Tramèr for their valuable guidance and ideas on this project.
@@ Acknowledgements
Appendix A Appendix
A.1 Base Audio File Details
@@
A.2 Judge Prompt Template
@@
A.3 Additional Results
@@
Experiment Original Drop Bandpass Recording Gaussian Denoise

### v2

@@ 3.1.3 Meaningfulness
While the adversarial noise alters model behavior, it is largely unintelligible to humans. This raises the question: how does the model interpret such input? The SALMON-N architecture allows inspection via two feature types: (1) BEATs labels, which provide discrete audio event predictions (e.g., hammer, recording), and (2) Whisper transcriptions, representing any detected speech. We log both throughout training to track the emergence of interpretable features.
@@ 3.2 Evaluation
We construct a set of 140 harmful prompts that the clean Vicuna model refuses to answer. These span seven categories (e.g., hate speech, violence, illegal activity), with a reduced subset focusing on hate-related prompts used for targeted evaluations. Prompt sources include prior red-teaming datasets (Qi et al., 2023; Gehman et al., 2020) and manual curation. We also include a 20-question control set from ARC-Easy (Clark et al., 2018) to assess performance on neutral tasks under jailbreak conditions.
For each jailbreak audio $x$, we prompt the model with harmful inputs $h_{i}$ and record $f(x,h_{i})$, truncating outputs to 150 tokens. Toxicity is assessed via the Detoxify API (Hanu & Unitary team, 2020) and Mixtral (Jiang et al., 2024) as an alignment judge. Details and judge prompts appear in Appendix  C.
@@ 4 Results
We run $178$ individual experiments and show that the resulting audio jailbreaks use the few-shot optimization corpus method, which exhibit a similar attack success rate (ASR) - albeit on a broader and different evaluation set - to visual jailbreaks generated by the same method (Qi et al., 2023).
Jailbreaks optimized on a few-shot corpus work, remain effective under stealth constraints, and can generalize across seemingly arbitrary dimensions.  Figure 1 shows up to 65% ASR on the specific target task from our few-shot optimization, and more powerfully, a considerable ASR on other misalignment tasks, showing generalization across a broad notion of “toxicity”. The jailbroken model shows low correlation (R² = 0.12) between logic accuracy and toxicity (Figure 4). Stealthy audio jailbreaks remain highly effective even under strong constraints, achieving up to 55% ASR with imperceptible perturbations (e.g., ϵitalic-$\epsilon\leq$ 0.001), as shown in Figure 3. Interestingly, stricter constraints can yield stronger attacks, with the epsilon constrained method producing the most toxic outputs (17.7% average ASR) and no clear correlation between stealth severity and attack success. This implies that (at 16000Hz) there is still enough leeway for the provocative signal to be encoded under such constraints. Finally, we show that a base-audio-agnostic 1s prepend snippet can be universally optimized to generalize across audio files, achieving 28.3% ASR on holdouts and up to 40% ASR without a suffix (Table 5). This universal snippet drives down the loss further than individually optimized audios, indicating stronger toxicity capture via multi-audio optimization (Figure 5), and evokes toxic output even from benign prompts - including no text and non-English inputs - which is is highly relevant to the stealthy single control scenario (Section 3).
Table 1: Results across different epochs for the music audio file, including Whisper Transcription and Beats Features.
@@
Audio jailbreaks succeed by hijacking the model’s linguistic priors, often masquerading as speech and adopting a toxic first-person voice. The model interprets adversarial noise as intelligible language, confirmed by coherent Whisper transcriptions and the consistent emergence of “Speech” in BEATs labels over time (Figures 2, 6 and 5). These transcriptions frequently begin with “I” and express explicit or sinister statements, suggesting the jailbreak crafts a toxic persona (e.g., “I’m a bigot man!” in Figure 2, “I’m not going to be mad…” in Figure 5). Spikes in ASR and toxicity often align with more offensive transcriptions (Figure 6). However, effective stealthy jailbreaks can sometimes show no such transcriptional clues - maintaining benign Whisper and BEATs outputs while still achieving 7–40% ASR - which shows that transcription-based signals are not a sufficient indicator of jailbreak success.
@@
Jailbreak success is hindered by signal degradation and accompanying prompts. Audio quality and filtering can reduce the Attack Success Rate (ASR). As shown in Table 3, over-the-air recording is the most damaging form of degradation, causing the largest average drop in ASR. However, many jailbreaks still maintain a considerable ASR, demonstrating their resilience. Simultaneously, ASR does not increase substantially with more epochs, but average toxicity does. Across the base audios, we observe that while more epochs generally lead to a gradual improvement in ASR, this increase is modest compared to the sharp rise in average toxicity throughout optimization (particularly evident in Figure 2 (left) and Figure 6). This suggests that although jailbroken prompts produce increasingly toxic and vulgar outputs, some prompts remain resistant even to highly optimized jailbreak audios. Characteristics of the base audio influence jailbreak optimization: frequency structure matters, with random noise enabling far stronger attacks than silence (Figure 7), suggesting some initial structure is necessary for optimization. Base audios with lower original loss tend to perform better under stealth constraints, likely due to requiring smaller perturbations (Figure 3), though no clear correlation is found with content, length, or transferability. Notably, even unoptimized random noise can occasionally trigger toxicity (8% ASR; Figure 4), underscoring the fragility of model alignment.
@@ 5 Conclusion and Future Work
We demonstrate that audio can be used to subvert language model alignment, with optimized audio often encoding toxic first-person speech. The success of minimally perturbed adversarial audio highlights the brittleness of current models and raises concerns for real-world systems, such as AI-powered robotics. Although we
Future Work. Our results open avenues for exploring jailbreaks across other audio-language models (Chu et al., 2023; Alayrac et al., 2022) and optimization objectives. Further work should study how different generation methods (Ying et al., 2024; Shayegani et al., 2023; Ma et al., 2024) and corpora affect the interpretability and generalization of jailbreaks, and how little information is needed (e.g., bits or $L_{\infty}$ perturbation) to encode an effective attack. Moreover, our interpretability results could inspire a similar investigation into image jailbreak features, and how the presence or absence of linguistic features might influence transferability (Schaeffer et al., 2024).
Defenses. Our findings show that transcription- or classification-based filters are insufficient: stealthy audio jailbreaks can evade detection while still producing harmful outputs. This has implications for both input sanitization and output filtering in audio-based systems.
@@
Appendix A Model Architecture Details
SALMON-N 7B is a multimodal audio-language model developed by Tsinghua University and ByteDance (Tang et al., 2024). It consumes audio by extracting two complementary sets of features:
  • BEATs features, which provide high-level audio event labels (e.g., Snicker, Drip, Human Sounds).
  • Whisper features, used primarily for speech transcription.
The audio signal is first converted into a spectrogram, from which these features are extracted. The extracted features are segmented into overlapping chunks and passed through a Q-former (Kim et al., 2024), which aligns them into a representation compatible with the language model input space.
The textual input is independently tokenized and embedded, and both the audio and text tokens are concatenated with a delimiter before being fed into the model. Cross-attention is applied over these mixed-modality input tokens.
The language modeling backbone is Vicuna-7Bv1.5 (Chiang et al., 2023), a model fine-tuned from LLaMA-7B on user-shared ChatGPT conversations. It inherits alignment via reinforcement learning from human feedback (RLHF) but lacks additional safety-specific filtering, making it susceptible to known jailbreak techniques (Chao et al., 2024).
We chose SALMON-N for its open-source availability, strong performance on tasks such as auditory question answering, emotion recognition, and audio captioning, as well as its well-structured combination of speech and non-speech features. Other ALMs such as Pengi (Deshmukh et al., 2024) and Qwen-Audio (Chu et al., 2023) adopt similar dual-extractor architectures.
Appendix B Base Audio File Details
@@
Appendix C Evaluation Details
C.1 Harmful Prompt Dataset
We evaluate jailbreak effectiveness using a curated set of 140 harmful prompts which the clean Vicuna-7Bv1.5 model reliably refuses to answer. These prompts span the following seven categories:
  • Target Hate: Derogatory content toward specific demographics or genders included in the jailbreak optimization dataset.
  • Non-Target Hate: Hate speech aimed at groups not present in the optimization data.
  • Hate Speech: Generally offensive prompts targeting individuals or humanity without group specificity.
  • Illegal Activity: Instructions or questions encouraging illegal behavior.
  • Physical Violence: Prompts glorifying or instructing violent actions.
  • Sexual Content: Sexually explicit or inappropriate prompts.
  • Other: Includes misinformation, political extremism, and doxxing requests.
Prompts were sourced from prior datasets such as Manual Harmful Instructions and Red Teaming Prompts (Qi et al., 2023), Real Toxicity Prompts (Gehman et al., 2020), and ChatGPT-generated samples, manually filtered for those to which the clean model refuses to respond.
For transferability experiments, we distinguish between the reduced evaluation set (only the first three hate-related categories) and the full evaluation set (all seven categories).
C.2 Control Set
To assess the unintended effects of jailbreaks on benign reasoning, we evaluate model accuracy on 20 multiple-choice logic and reasoning questions sampled from the ARC-Easy dataset (Clark et al., 2018). We report accuracy as $logic(f,x)$ under jailbreak audio input $x$.
C.3 Toxicity Judging Procedure
We generate model responses $f(x,h_{i})$ for each harmful prompt $h_{i}$ with audio input $x$, limiting outputs to 150 tokens. Toxicity is evaluated in three ways:
  1. 1.
    Detoxify API (Hanu & Unitary team, 2020): Predicts toxicity scores from 0 to 1 for free-text outputs.
  2. 2.
    Mixtral-8x7B-Instruct Judge (Jiang et al., 2024): A language model judge prompted to classify outputs as toxic or not based on alignment categories of interest.
  3. 3.
    Manual Audit: Human reviewers check a random sample of outputs to validate model and API judgments.
The prompt template used for the Mixtral judge is provided in Section C.4.
C.4 Judge Prompt Template
@@
Appendix D Jailbreak Generation Details
D.1 Optimization Objective
We follow Qi et al. (2023) to optimize adversarial audio inputs. Given a base audio $x_{0}$, target corpus $t=\{t_{0},\dots,t_{n}\}$, and a differentiable model $f$, we minimize the cross-entropy loss between the model’s output distribution and the target text:
$x_{\text{adv}}=\arg\min_{x}-\sum_{i=0}^{n}t_{i}\log P_{f}(t_{i}|x)$
This yields audio $x_{\text{adv}}$ that increases the likelihood of generating the harmful targets $t$ when paired with any prompt. During training, we use empty textual input to enforce prompt-agnosticity. The target corpus consists of 66 manually curated toxic sentences spanning identity groups and generalized hate, as in (Qi et al., 2023). We randomly select 8 target sentences per epoch.
D.2 Stealth Constraints
To simulate more realistic or covert scenarios, we apply three types of stealth constraints:
Epsilon-Constrained Optimization
We bound the maximum allowed change to each sample in $x$ by clipping the update within a margin $\epsilon$:
$\forall i:x_{t+1}[i]=\text{clip}(x_{t}[i]-\eta\nabla_{x}\mathcal{L}(x_{t},t),\,x_{0}[i]-\epsilon,\,x_{0}[i]+\epsilon)$
We evaluate $\epsilon\in\{0.1,0.01,0.001,0.0001\}$ to explore trade-offs between stealth and effectiveness.
Frequency-Hiding (Band-Stop Filtering)
We hide adversarial perturbations in specific frequency bands of a perturbed audio $x$, removing components in the range $[b_{l},b_{u}]$ via:
$\hat{x}[f]=\begin{cases}x[f],&\text{if }f<b_{l}\text{ or }f>b_{u}\\ 0,&\text{if }b_{l}\leq f\leq b_{u}\end{cases}$
Here $x[f]$ denotes the Fourier-transformed audio at frequency $f$. We experiment with frequency bands: $(1000,8000)$, $(100,10000)$, $(40,20000)$, and $(50,15000)$.
Prepend Optimization
Instead of modifying the original audio, we optimize a short prefix $p$ of length $d$ seconds:
$p^{*}=\arg\min_{p\in[-1,1]^{16000d}}\mathcal{L}([p\|x],t)$
We experiment with $d\in\{2,1,0.1,0.01\}$. The prefix is initialized with random noise in $[-1,1]$.
D.3 Audio-Agnostic Generalization
To achieve prompt- and audio-agnostic jailbreaks, we train a single prefix $p$ jointly across a set of base audios $B=\{x^{1},\dots,x^{n}\}$:
$\mathcal{L}_{\text{total}}=\frac{1}{|B|}\sum_{x\in B}\mathcal{L}([p\|x],t)$
We evaluate generalization by holding out one base audio for testing after training on the remaining $n-1$. We use $d=1$ second and clip sample amplitudes of $p$ to ±$\pm 0.1$, making it perceptible as a brief noise burst. This setup ensures the resulting adversarial prefix is robust across different base audio contexts.
Appendix E Additional Results
@@
Experiment Original Intermittent Silence Masking Bandpass Filtering Over the Air Recording Gaussian Noise Removal
@@
Epochs Whisper Beats
Holdout Audio ASR (%) Avg Toxicity
Music 25.0% 0.088
Mountain 23.3% 0.082
Gunshots 36.7% 0.204
Duck 36.7% 0.190
Excitement 20.0% 0.136
Jailbreak Snippet Only 40.0% 0.240